bralyn:

What is the solution to this equation? 2 x + 10 = 28

KazGotFetee:

To find what X equals?

FredrickStar:

do yk how to factor?

KazGotFetee:

Well to find what x Equals we See the Question 2x + 10 = 28 you would subtract 10 to each sides 10 and 28 what would you get?

KazGotFetee:

so 10-10 and 28-10 = what?

FredrickStar:

10-10 = 0 and 28-10=18

KazGotFetee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @FredrickStar 10-10 = 0 and 28-10=18 \(\color{#0cbb34}{\text{End of Quote}}\) ok good now divide 2x by 2 and 18 by 2 what do we get.

FredrickStar:

mhm

FredrickStar:

18 divided by 2 and =9 which means x=9

KazGotFetee:

Correct!!!

FredrickStar:

mhm

FredrickStar:

good job

KazGotFetee:

@bralyn

KazGotFetee:

there's your answer!

KazGotFetee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @FredrickStar good job \(\color{#0cbb34}{\text{End of Quote}}\) you too!!!